# Reply with Sticker: New Dataset and Model for Sticker Retrieval

### v3

@@ Abstract
Using stickers in online chatting is very prevalent on social media platforms, where the stickers used in the conversation can express someone’s intention/emotion/attitude in a vivid, tactful, and intuitive way. Existing sticker retrieval research typically retrieves stickers based on context and the current utterance delivered by the user. That is, the stickers serve as a supplement to the current utterance. However, in the real-world scenario, using stickers to express what we want to say rather than as a supplement to our words only is also important. Therefore, in this paper, we create a new dataset for sticker retrieval in conversation, called StickerInt, where stickers are used to reply to previous conversations or supplement our words [^1] We believe that the release of this dataset will provide a more complete paradigm than existing work for the research of sticker retrieval in the open-domain online conversation.. Based on the created dataset, we present a simple yet effective framework for sticker retrieval in conversation based on the learning of intention and the cross-modal relationships between conversation context and stickers, coined as Int-RA. Specifically, we first devise a knowledge-enhanced intention predictor to introduce the intention information into the conversation representations. Subsequently, a relation-aware sticker selector is devised to retrieve the response sticker via cross-modal relationships. Extensive experiments on the created dataset show that the proposed model achieves state-of-the-art performance in sticker retrieval [^2] The dataset and source code of this work are released at https://github.com/HITSZ-HLT/Int-RA..
@@ I Introduction
Recent research endeavors of sticker retrieval have been dedicated to using stickers to supplement the current response, in order to strengthen the expression of emotion, attitude, or opinion [7, 8]. However, in the real-world scenario, we may also use stickers to reply to the previous conversation directly, rather than merely supplementing our words with stickers. One expected scenario of the sticker retrieval task is that suitable stickers can be retrieved for replies whether or not one has made a textual response. To illustrate our idea, we give examples shown in Figure 1. Figure 1 (a) shows the sticker retrieval scenario of using a sticker to supplement one’s current utterance. Figure 1 (b) is another sticker retrieval scenario, in which the user uses a sticker to reply to previous conversations directly. Therefore, in our work, we create a new sticker retrieval dataset to cover these two scenarios, called StickerInt, which is a comprehensive consideration for using stickers in social media conversation. Our StickerInt dataset contains 1,578 Chinese conversations with 12,644 utterances.
@@
  • Experiments conducted on our StickerInt dataset demonstrate that our proposed framework outperforms the baseline models.
@@ II-B Image Retrieval
Image retrieval aims to retrieve images that match a given query from a large collection of images. Early method [11] mainly relied on global feature-based image-text matching. Further, [12] proposed a Visual Semantic Reasoning Network (VSRN), which generates global features with regional semantic associations through a Graph Convolutional Network (GCN). To address the false negative problem in existing image-text matching benchmarks, [13] proposed an ITM framework integrating Language Guidance. In terms of improving retrieval efficiency and accuracy, [14] proposed a self-supervised fine-grained alignment module called SelfAlign. [15] proposed a feature-first approach to advance image-text retrieval by improving visual features. This method generates more comprehensive and robust visual features through a two-step interaction model and a multi-attention pooling module. Recently, with the development of large-scale pretraining models like CLIP [16] and ALIGN [17], significant performance improvements are achieved in image-text matching tasks by pretraining on massive vision-language data and then fine-tuning for specific downstream tasks. Different from the image retrieval methods, we focus on retrieving the image based on the conversation not just the text.
@@ II-C Multi-modal Conversation
Several multi-modal studies have sought to improve the efficacy of conversational agents by enriching textual expressions of generated dialog responses through associative vision scenes [18, 19, 20, 21]. In contrast, sticker retrieval in the real-world scenario requires understanding the semantic expression and intention of context and the sticker image. Several works specifically paid attention to sticker features [22, 4]. [22] proposed a real-time system for sticker recommendation, which decomposes the sticker recommendation task into two steps including message prediction that the user is likely to send and an appropriate sticker substitution. Nowadays, using stickers for replies has become commonplace in social media interactions. There has been a growing number of works on sticker retrieval, which assists users in selecting the appropriate sticker for response. [8] proposed a sticker response selector model that employs a convolutional-based sticker image encoder and a self-attention-based multi-turn dialog encoder to obtain the representation of stickers and utterances. Then a deep interaction network is designed for obtaining deep matching and a fusion network is employed to output the final matching score. [7] presented a Meme incorporated Open-domain Dialogue (MOD) task and utilized pooling all sub-tasks like text generation and internet meme prediction into a unified sequence generation procedure to solve it. To tackle three challenges of inherent multimodality, significant inter-series variations, and complex multi-modal sentiment fusion. [4] integrated both the sticker text and series tag to holistically model sticker sentiment. They employed a flexible masked attention mechanism to selectively extract the most pertinent information crucial for the current sentiment assessment. These methods which directly match conversation context and stickers, overlook the expressive role of stickers in conversation, where emotions and intentions are conveyed in a visually engaging manner.
@@ III StickerInt Dataset
This section introduces our new dataset StickerInt for sticker retrieval in detail. Specifically, Section III-A presents the construction of our StickerInt dataset, Section III-B introduces the annotation process of StickerInt, and Section III-C analyzes the created StickerInt by data statistics.
@@ III-A Data Construction
We collected our dataset from a widely used social platform (WeChat [^3] https://weixin.qq.com/), which boasts a vast of conversations with stickers accessible for both individual and group chats. We select four open chat groups with active participants and collect their conversations. Among them, each group engages in open-domain online conversation, making the use of stickers more diverse. Note that we eliminate extraneous image elements like screenshots and photos.
We formulate stringent guidelines and policies for data preprocessing. To protect user privacy, users’ personal information (including real name, age, address, etc.) is deleted, and user IDs are anonymized in the data. Additionally, we exclude any content that may contain inappropriate, offensive, or insulting expressions. Furthermore, we segment the whole chat content within the dataset into multiple conversations to ensure the integrality and independence of each conversation. Based on this, we traverse each sticker in the chat history, capturing its associated context to derive conversations with stickers, which ensures that each sticker has a corresponding conversation context.
TABLE I: Statistics of StickerInt dataset. con. = conversation, Avg. = average.
@@ III-B Data Annotation
Figure 3: Similarity distribution among all stickers in our dataset. The x-axis represents the count of SSIM calculation, and the y-axis indicates the different range of sticker similarity.
@@ III-C Dataset Analysis
The detailed statistics of our dataset are shown in Table I. In total, there are 1,578 conversations which contain 12,644 utterances, 59,623 tokens, and 1,025 stickers. Each conversation includes 8.01 utterances on average. The average number of users who participate in a conversation is 2.80. The average number of tokens in an utterance is 5.21. Furthermore, we also visualize the distribution of intention tags in Figure 2. We can see that the proportion of stickers used in different labels varies, which also demonstrates the diversity of stickers’ intention expressions in our dataset.
@@ IV Methodology
In this section, we introduce our proposed Int-RA framework for sticker retrieval in detail. Assume that there is a conversation context $T=\{t_{1},...,t_{N_{t}},s_{j}\}$ and a sticker set $V=\{v_{1},...v_{N_{v}}\}$, where $t_{i}=(s_{i},u_{i})$, $s_{i}$ and $u_{i}$ represent the $i$-th user and utterance. $N_{t}$ and $N_{v}$ represent the number of utterances and stickers, respectively. The sticker retrieval task aims to select a suitable sticker from the sticker set $V$ based on the conversation context $T$ for the user $s_{j}$. In the $i$-th utterance $u_{i}=\{w_{1}^{i},...,w^{i}_{N_{w}^{i}}\}$, $w_{j}^{i}$ represents the $j$-th word in $u_{i}$, and $N_{x}^{i}$ represents the total number of words in $u_{i}$ utterance. Additionally, our dataset presents an annotated intention tag $y_{int}$ as supplemental information in each conversation $T$. In the dialog context $T$, $s_{i}$ represents a sticker image with a binary label $y_{i}$, indicating whether $s_{i}$ is an appropriate response for $T$. Each sticker with the intention label $m_{i}$ indicates the intention of the speaker to use the sticker to respond. Our goal is to learn a ranking model that can produce the correct ranking for each dialog.
@@ IV-B Attribute-aware Sticker Representation
To extract the key expression information and reduce unnecessary interference from irrelevant information, for the learning of stickers, we first devise four visual attributes, i.e. gesture $L_{G}$, posture $L_{P}$, facial expression $L_{F}$, and verbal $L_{V}$, to construct prompts for the strikers. Based on this, we use Qwen-VL [29] as the MLLM to produce attribute-aware sticker descriptions based on the above four attributes:
@@ IV-C Relation-aware Sticker Selector
where $\rho_{neg}$ and $\rho_{pos}$ correspond to the cosine similarity of non-true (negative) stickers and true (positive) stickers. The margin is the margin rescaling.
@@ V Experiment
This section details the experimental settings and experimental results of our proposed Int-RA framework conducted on the created StickerInt dataset. We first present the experimental settings in Section V-A. Then, we introduce various compared methods in Section V-B. We analyze the performance of our approach through the Main Results (V-C), Ablation Study (V-D), Effect of Number of Utterances (V-F), and Effect of Different Attributes (V-G). Finally, to more intuitively demonstrate the performance of our method, we visualize the relations in Section V-I and present some interactive cases in Section V-H.
@@ V-A Experimental Settings
Implement details. We adopt Multi-lingual BERT [28] as the text encoder to derive the textual representation. The CLIP pre-trained ViT model [30] is employed as the image encoder to derive the visual representation. we adopt Qwen-VL [29] to generate sticker descriptions  [^4] In the preliminary experiment, we also try other MLLMs such as mini-GPT4 [31] and Llava [32]. We found that Qwen-VL performed slightly better.. We set the batch size to 4 and use Adam optimizer [33] as our optimizing algorithm. The learning rate is set to $1\times 10^{-4}$. Both $\lambda_{1}$ and $\lambda_{2}$ are set to 1.
Evaluation metrics. Three widely used evaluation metrics are applied in our experiments: mean of average precision (mAP), top N-precision (P@N).mAP is a widely accepted criterion for assessing retrieval accuracy [34]. P@N evaluates the precision of the top N predictions. Here, we mainly present the results for P@1, P@3, and P@5. Moreover, the PR curve visually illustrates the trade-off between precision and recall at various thresholds. Notably, if the retrieved sticker matches the intention label of the ground truth sticker, we consider the result correct, as multiple stickers can serve as responses to the same conversation.
@@ V-B Compared Methods
To evaluate the performance of our model, we compare the proposed Int-RA with several baseline methods, including existing sticker retrieval methods, recent text-to-image retrieval approaches, and large language models (LLMs).
  • Sticker retrieval methods: MOD [35], which leverages a unified generation network to produce multi-modal responses. SRS [36] which learns sticker representations and utterance context in the multi-turn dialog.
  • Text-to-image retrieval methods: LGUR [37] which enhances feature granularity alignment and performance through a transformer-based approach. IRRA [38] which leverages implicit relation reasoning and similarity distribution matching for improved cross-modal alignment. PCME [39] which presents a probabilistic embedding model for cross-modal retrieval. CLIP [16] which introduces a vision model trained on internet image-text pairs.
  • Large language models: Baichuan2, LLama3, ChatGLM3, Qwen2, Qwen-VL, and LLaVA.
@@
Int-RA (ours) 18.18 37.66 40.91 53.37

### v4

@@ Abstract
Using stickers in online chatting is very prevalent on social media platforms, where the stickers used in the conversation can express someone’s intention/emotion/attitude in a vivid, tactful, and intuitive way. Existing sticker retrieval research typically retrieves stickers based on context and the current utterance delivered by the user. That is, the stickers serve as a supplement to the current utterance. However, in the real-world scenario, using stickers to express what we want to say rather than as a supplement to our words only is also important. Therefore, in this paper, we create a new dataset for sticker retrieval in conversation, called StickerInt, where stickers are used to reply to previous conversations or supplement our words. Based on the created dataset, we present a simple yet effective framework for sticker retrieval in conversation based on the learning of intention and the cross-modal relationships between conversation context and stickers, coined as Int-RA. Specifically, we first devise a knowledge-enhanced intention predictor to introduce the intention information into the conversation representations. Subsequently, a relation-aware sticker selector is devised to retrieve the response sticker via cross-modal relationships. Extensive experiments on two datasets show that the proposed model achieves state-of-the-art performance and generalization capability in sticker retrieval [^1] The dataset and source code of this work are released at https://github.com/HITSZ-HLT/Int-RA..
@@ I Introduction
Recent research endeavors of sticker retrieval have been dedicated to using stickers to supplement the current response, in order to strengthen the expression of emotion, attitude, or opinion [7, 8]. However, in the real-world scenario, we may also use stickers to reply to the previous conversation directly, rather than merely supplementing our words with stickers. We refer to the former as the Supplementary Response (SR) scenario, where stickers are used alongside the speaker’s own utterance, and to the latter as the Direct Response (DR) scenario, where stickers function as independent replies without accompanying text. One expected scenario of the sticker retrieval task is that suitable stickers can be retrieved for replies whether or not one has made a textual response. To illustrate our idea, we give examples shown in Figure 1. Figure 1 (a) shows the SR scenario, and Figure 1 (b) depicts the DR scenario. Therefore, in our work, we create a new sticker retrieval dataset to cover these two scenarios, called StickerInt, which is a comprehensive consideration for using stickers in social media conversation. Our StickerInt dataset contains 1,578 Chinese conversations with 11,066 utterances.
@@
  • Experiments conducted both on our created StickerInt and a public dataset demonstrate that our proposed framework outperforms the baseline models.
@@ II-B Image Retrieval
Image retrieval aims to retrieve images that match a given query from a large collection of images. Early method [11] mainly relied on global feature-based image-text matching. Further, [12] proposed a Visual Semantic Reasoning Network (VSRN), which generates global features with regional semantic associations through a Graph Convolutional Network (GCN). To address the false negative problem in existing image-text matching benchmarks, [13] proposed an Image-Text Matching (ITM) framework integrating Language Guidance. In terms of improving retrieval efficiency and accuracy, [14] proposed a self-supervised fine-grained alignment module called SelfAlign. [15] proposed a feature-first approach to advance image-text retrieval by improving visual features. This method generates more comprehensive and robust visual features through a two-step interaction model and a multi-attention pooling module. Recently, with the development of large-scale pretraining models like CLIP [16] and ALIGN [17], significant performance improvements have been achieved in ITM tasks by pretraining on massive vision-language data and then fine-tuning for specific tasks. Different from the image retrieval methods, we focus on retrieving the image based on the conversation not just the text.
@@ II-C Multi-modal Conversation
Several multi-modal studies have sought to improve the efficacy of conversational agents by enriching textual expressions of generated dialog responses through associative vision scenes [18, 19]. In contrast, sticker retrieval in the real-world requires understanding the semantic expression and intention of the context and the sticker image. Several works specifically paid attention to sticker features [20, 4]. [20] proposed a real-time system for sticker recommendation, which decomposes the sticker recommendation task into two steps, including message prediction that the user is likely to send and an appropriate sticker substitution. Nowadays, using stickers for replies has become commonplace in social media interactions. There has been a growing number of works on sticker retrieval, which assists users in selecting the appropriate sticker for a response. [8] proposed a sticker response selector model that employs a convolution-based sticker image encoder and a self-attention-based multi-turn dialog encoder to obtain the representation of stickers and utterances. Then, a deep interaction network is designed for obtaining deep matching, and a fusion network is employed to output the final matching score. [7] presented a Meme incorporated Open-domain Dialogue (MOD) task and utilized pooling all sub-tasks like text generation and internet meme prediction into a unified sequence generation procedure to solve it. To tackle three challenges of inherent multimodality, significant inter-series variations, and complex multi-modal sentiment fusion, [4] integrated both the sticker text and series tag to holistically model sticker sentiment. They employed a flexible masked attention mechanism to selectively extract the most pertinent information crucial for the current sentiment assessment. These methods, which directly match conversation context and stickers, overlook the expressive role of stickers in conversation, where emotions and intentions are conveyed in a visually engaging manner.
@@ III StickerInt Dataset
This section introduces our new dataset StickerInt for sticker retrieval in detail. Specifically, Section III-A presents the construction of our StickerInt dataset, Section III-B introduces the annotation process of StickerInt, and Section III-C and Section III-D analyze the created StickerInt dataset.
@@ III-A Data Construction
We collected our dataset from a widely used social platform (WeChat [^2] https://weixin.qq.com/), which boasts a vast number of conversations with stickers accessible for both individual and group chats. We select four open chat groups with active participants and collect their conversations. Among them, each group engages in open-domain online conversation, making the use of stickers more diverse. Note that we eliminate extraneous image elements like screenshots and photos.
We formulate stringent guidelines and policies for data preprocessing. Specifically, we follow established best practices adopted in prior work such as [21] and [22], and implement a two-stage anonymization pipeline. (1) In the first stage, we conduct automated removal and replacement of direct identifiers. All direct user-related information, including user IDs, phone numbers, nicknames, and profile metadata, is systematically removed or replaced with pseudonyms. To preserve the continuity and coherence of multi-turn conversations, we assign consistent anonymized tokens such as "User_1" and "User_2" to represent different participants. Usernames are replaced rather than deleted to retain clarity in speaker attribution while safeguarding privacy. (2) In the second stage, we perform a manual review to identify and anonymize any remaining indirect identifiers. These may include references to personal occupations, geographic locations, interpersonal relationships, or infrequent linguistic patterns that could potentially lead to re-identification. For such instances, we apply additional anonymization strategies such as data generalization to reduce specificity and data suppression to remove sensitive elements entirely.
In addition to protecting user privacy, we remove any content that contains inappropriate, offensive, or abusive language to comply with ethical research standards. To ensure contextual integrity, we segment the chat history into distinct and coherent conversations. Each sticker in the dataset is aligned with its surrounding conversational context, guaranteeing that it is situated within a meaningful and self-contained dialogue.
TABLE I: Statistics of StickerInt dataset. con. = conversation, Avg. = average. SR and DR represent Supplementary Response and Direct Response Scenarios.
@@ III-B Data Annotation
Figure 3: Similarity distribution among all stickers in our dataset. The x-axis represents different ranges of sticker similarity, and the y-axis indicates the count of SSIM calculations.
@@ III-C Dataset Analysis
The detailed statistics of our dataset are shown in Table I. In total, there are 1,578 conversations that contain 11,066 utterances, 59,623 tokens, and 1,025 stickers. Of these stickers, 702 are associated with textual content. The proposed dataset is collected from conversations with Chinese users on WeChat, and the dialogues are predominantly in Chinese. Notably, 841 utterances include English words such as “OK”, “QQ”, and “NLP”, which are naturally embedded in Chinese expressions, reflecting the way people commonly communicate in daily life. Each conversation includes 8.01 utterances on average. The average number of users who participate in a conversation is 2.80. Among all conversations in our dataset, 1,017 belong to the Supplementary Response scenario, and 561 fall under the Direct Response scenario (sticker as a direct reply to previous conversation turns). The average number of tokens in an utterance is 5.21.
Furthermore, we also visualize the distribution of intention tags in Figure 2. We can see that the proportion of stickers used in different labels varies, which also demonstrates the diversity of stickers’ intention expressions in our dataset.
@@ IV Methodology
In this section, we introduce our proposed Int-RA framework for sticker retrieval in detail. Assume that there is a conversation context $T=\{t_{1},...,t_{N_{t}},s_{j}\}$ and a sticker set $V=\{v_{1},...v_{N_{v}}\}$, where $t_{i}=(s_{i},u_{i})$, $s_{i}$ and $u_{i}$ represent the $i$-th user and utterance. $N_{t}$ and $N_{v}$ represent the number of utterances and stickers, respectively. The sticker retrieval task aims to select a suitable sticker from the sticker set $V$ based on the conversation context $T$ for the user $s_{j}$. In the $i$-th utterance $u_{i}=\{w_{1}^{i},...,w^{i}_{N_{w}^{i}}\}$, $w_{j}^{i}$ represents the $j$-th word in $u_{i}$, and $N_{w}^{i}$ represents the total number of words in $u_{i}$ utterance. Additionally, our dataset presents an annotated intention tag $y_{int}$ as supplemental information in each conversation $T$. In the dialog context $T$, $v_{i}$ represents a sticker image with a binary label $y_{i}$, indicating whether $s_{i}$ is an appropriate response for $T$. Each sticker with the intention label $m_{i}$ indicates the intention of the speaker to use the sticker to respond. Our goal is to learn a ranking model that can produce the correct ranking for each dialog.
@@ IV-B Attribute-aware Sticker Representation
To capture key expressive features and reduce interference from irrelevant style variations in sticker learning, we adopt four representative attributes informed by prior research in visual communication and multimodal semantics. Studies have shown that gesture, facial expression, posture, and verbal elements are fundamental to the visual encoding of meaning. [29] highlights these visual signs as core denotative elements whose interpretation depends on contextual knowledge. [30] further demonstrates their frequent use in memes to convey intention and emotion. Guided by these findings, we define four attributes for sticker representation: gesture $L_{G}$, posture $L_{P}$, facial expression $L_{F}$, and verbal text $L_{V}$. Together, these features encompass both visual and textual cues, enabling more robust and fine-grained semantic modeling for sticker retrieval. Based on this, we use Qwen-VL [31] as the MLLM to produce attribute-aware sticker descriptions based on the above four attributes:
@@ IV-C Relation-aware Sticker Selector
where $\rho_{neg}$ and $\rho_{pos}$ correspond to the cosine similarity of non-true (negative) stickers and true (positive) stickers. The margin is the rescaling margin.
@@ V Experiment
This section details the experimental settings and experimental results of our proposed Int-RA framework conducted on the created StickerInt dataset. We first present the experimental settings in Section V-A. Then, we introduce various compared methods in Section V-B. We analyze the performance of our approach through the Main Results (V-C), Results on Other Datasets (V-D), Ablation Study (V-E), Effect of Using Intention (V-F), Effect of Number of Utterances (V-G), and Effect of Different Attributes (V-H). Subsequently, we delve into the computational efficiency and scalability in Section V-I. Finally, to more intuitively demonstrate the performance of our method, we present some interactive cases in Section V-J and visualize the relations in Section V-K.
@@ V-A Experimental Settings
Implement details. We adopt Multi-lingual BERT [28] as the text encoder to derive the textual representation. The CLIP pre-trained ViT model [32] is employed as the image encoder to derive the visual representation. We adopt Qwen-VL [31] to generate sticker descriptions [^3] In the preliminary experiment, we also try other MLLMs such as mini-GPT4 [33] and Llava [34]. We found that Qwen-VL performed slightly better.. We set the batch size to 4 and use Adam optimizer [35] as our optimization algorithm. The learning rate is set to $1\times 10^{-4}$. Both $\lambda_{1}$ and $\lambda_{2}$ are set to 1.
Dataset. To evaluate the generalization ability of our method, we additionally conduct experiments on MOD [7] dataset, which includes bilingual (Chinese and English) dialogues and supports several tasks; we focus on the meme retrieval task. Each conversation has an average of 13.42 turns and 11.46 tokens per turn. As the official test set is unavailable, we report results on the validation set.
Evaluation metrics. Two widely used evaluation metrics are applied in our experiments: mean of average precision (mAP), top N-precision (P@N). mAP is a widely accepted criterion for assessing retrieval accuracy [36]. P@N evaluates the precision of the top N predictions. Here, we mainly present the results for P@1, P@3, and P@5. Notably, if the retrieved sticker matches the intention label of the ground truth sticker, we consider the result correct, as multiple stickers can serve as responses to the same conversation. For the MOD dataset, we follow the original setting using recall at position k in n candidates Rn@k as an evaluation metric, which measures if the positive response is ranked in the top k positions of $n$ candidates.
@@ V-B Compared Methods
To evaluate the performance of our model, we compare the proposed Int-RA with several baseline methods, including existing sticker retrieval methods, recent text-to-image retrieval approaches, and large language models-based methods.
  • Sticker retrieval methods: MOD [37], which leverages a unified generation network to produce multi-modal responses. SRS [21], which learns sticker representations and utterance context in the multi-turn dialog. For a fair comparison, these two models are fine-tuned on our StickerInt dataset.
  • Text-to-image retrieval methods: LGUR [38] enhances feature granularity alignment via a transformer-based framework. IRRA [39] improves cross-modal alignment through implicit relation reasoning and similarity distribution matching. PCME [40] introduces a probabilistic embedding model for cross-modal retrieval. CLIP, Chinese-CLIP [16] and ViLT [41] are encoder-based vision-language models, fine-tuned for sticker selection. BLIP-2 [42] and FROMAGe [43] are encoder-decoder models. Given that FROMAGe is built on a large language model and pre-trained on large-scale multimodal dialogue data, we apply it directly to the task without additional fine-tuning.
  • Large language models-based methods: Baichuan2, Llama3, ChatGLM3, Qwen2, Qwen-VL, and LLaVA. Thanos [44] infuses skill-of-mind reasoning into LLMs, generating both responses and conversational skills. Ultron [45] is a multimodal conversational agent that incorporates social and contextual dynamics for image retrieval. ChatIR [46] is a dialogue-driven image retrieval system that interacts with users to iteratively refine queries and retrieve target images.
@@
-SR scenario 20.78 39.61 41.56 56.79
-DR scenario 13.64 28.57 38.31 54.75
V-C Main Results
For text-to-image retrieval methods, they prioritize capturing semantic relationships between text and image content. However, since they are not explicitly designed for sticker retrieval scenarios, they exhibited inferior performance compared to our framework. In addition, an interesting observation is the overall superior performance of text-to-image retrieval methods compared to sticker retrieval baseline methods. This disparity can be attributed in part to the model design of SRS and MOD, where both models are devised to retrieve suitable stickers from a limited set of similar sticker candidates. Consequently, more attention is devoted to distinguishing local information among similar sticker expressions. In contrast, our dataset encompasses a more diverse range of stickers in real-world conversations, requiring the model to analyze and recognize more complex visual features. This demonstrates the advantage of the new dataset and the proposed framework in this work.
For large language models, except for Ultron and ChatIR, we design a prompt that integrates the current session to generate the sticker description for each session, leveraging the in-context learning ability of LLMs [47, 48]. We then retrieve the appropriate sticker based on the generated sticker description and the combination of intent and sticker attributes using OpenAI’s LLM-based embedding model (text-embedding-ada-003) [49]. We can observe that LLM models perform better than sticker retrieval and text-to-image retrieval methods. This superior performance is attributed to the extensive parameterization and complex network architecture of LLMs, which greatly enhance their ability to understand and generate complex language and image descriptions. For P@1, Ultron performs the best, while for P@3 and P@5, ChatGLM3 achieves the highest performance, reaching 25.32% and 31.17%. Compared to our Int-RA method, all baseline models perform significantly worse. This further emphasizes the effectiveness of our approach in intent derivation, highlighting the critical role of intent as a bridging component in the process.
Results of different scenarios. As illustrated in Figure 1, our sticker retrieval task covers two typical usage scenarios: SR scenario, where a sticker is used to supplement the current utterance, and DR scenario, where a sticker is used to directly reply to a previous conversation without accompanying text. To assess the performance of our model in these two real-world settings, we conduct separate evaluations, with results summarized in Table V-B. The comparison reveals that the model performs better under the SR scenario than the DR scenario, suggesting that retrieval is more effective when the sticker is used to complement explicit textual input. In contrast, the DR scenario presents greater challenges, as it requires the model to infer implicit intent from the broader dialogue context without direct linguistic cues. This discrepancy highlights the increased complexity of sticker retrieval in context-dependent, intention-driven interactions, and further emphasizes the necessity of intention modeling and commonsense reasoning in such cases.
TABLE III: Experimental results (%) of various methods in MOD dataset. Bold indicates the best results, and underline denotes the second-best. represents the results from [50].
Model R10@1 R10@3 R10@5
FROMAGe 9.83 28.96 49.85
BLIP-2 17.64 35.29 47.18
ViLT 17.31 40.00 63.27
CLIP 21.90 47.72 66.04
MOD 21.93 41.33 70.78
Int-RA (ours) 21.90 44.50 70.90
V-D Results on MOD Dataset
We evaluate our proposed method Int-RA on a public benchmark dataset, MOD, to examine its generalization capability beyond our own StickerInt dataset. As shown in Table III, Int-RA achieves a R10 @ 5 of 70. 90%, slightly surpassing the best reported result by MOD. These results indicate that Int-RA, while not always the top-performing model across all metrics, can achieve performance on par with the best existing methods across diverse datasets, demonstrating strong generalization and robustness.
TABLE IV: Results of ablation study (%). w/o means without.
Model P@1 P@3 P@5 mAP
Int-RA 18.18 37.66 40.91 53.37